# A Multi-Fidelity Parametric Framework for Reduced-Order Modeling using Optimal Transport-based Interpolation: Applications to Diffused-Interface Two-Phase Flows

### 04 Mar 2026 v1

@@ 6.2. Rayleigh-Taylor instability
As an additional test case, we consider the simulation of a Rayleigh-Taylor instability, which occurs when an interface between two fluids with different densities experiences a pressure gradient opposing the density gradient. The computational domain is $d\times 4d]$ with $d=1$, and the interface is initially defined as the curve $y(x)=2d+0.1d\cos{(2\pi x/d)}$. The Rayleigh-Taylor instability is characterized by the Reynolds number $\mathrm{Re}=(\rho_{1}d^{3/2}\|\textbf{g}\|^{1/2})/\mu$ and the Atwood number $\mathrm{At}=(\rho_{1}-\rho_{2})/(\rho_{1}+\rho_{2})$ which are classically set to $3000$ and $0.5$, time is adimensionalized as $t^{*}=t/\sqrt{d/(\|\textbf{g}\|\mathrm{At})}$, while for the EOS relation in Equation (27) we chose the following parameters $\gamma_{1}=\gamma_{2}=1.4$, $P^{\infty}_{1}=P^{\infty}_{2}=6000$.
The Reynolds number characterizes the relative importance of inertial to viscous effects in the coupled system and is governed by gravity and the mixture viscosity, which is assumed constant across the two phases for this case. The Atwood number, on the other hand, quantifies the density contrast between the phases. When $\mathrm{At}=0$, the two phases have identical densities, whereas increasing values of the Atwood number correspond to progressively larger density ratios. As the Atwood number increases, one therefore expects the flow to develop increasingly complex interfacial patterns.
Concerning its parametric extension, we considered varying configurations corresponding to Atwood numbers in the following range:
@@
An overview of the parametric problem’s dynamics is shown in Figure 19.
@@
Figure 19. Phase-field isocontours at different Atwood numbers for the Rayleigh-Taylor instability at $t^{*}=3.0$. First row: low-fidelity simulations. Second row: high-fidelity simulations.
The top boundary is treated as a Riemann-invariant boundary condition with zero velocity and constant pressure, the bottom boundary is a no-slip wall whereas slip-wall boundary conditions are prescribed on the other lateral sides of the domain. For this study, two different meshes with quadrilateral elements were considered, involving 4040k and 160160k degrees of freedom. A $4^{\mathrm{th}}$ order Runge-Kutta scheme was used for time integration. If not specified otherwise, we will always consider the $\mathrm{At}=0.35$ case to test our framework for the Rayleigh-Taylor Instability.
@@ 6.2.1. MF-OT-ROM
A markedly different trend emerges compared to the previous test case: here, the dynamics accelerates over time. Initially, the interface evolves relatively slowly, but it subsequently accelerates due to gravitational effects, leading to complex breakup patterns. This behavior is also reflected in the MF-OT-ROM procedure, as the errors tend to grow as time progresses. This expresses that a more involved regression of the virtual time $\alpha$ may be needed to better capture the accelerating dynamics, but the current approximation already provides a significant improvement over the uncorrected low-fidelity solution, even with a limited amount of data, and increasing the number of checkpoints significantly improves the overall accuracy of the method, consistently with the behavior observed in the previous test case.
@@
In Figure 21, we present the full predicted field of $\phi_{1}$ for the Rayleigh-Taylor instability problem. Each panel is split into two halves, where the high-fidelity solution is mirrored against the corresponding low-fidelity prediction (either corrected or uncorrected) to facilitate visual comparison. From left to right, the level of OT correction increases. The left panel displays the uncorrected low-fidelity solution, while the central and right panels show the corrected low-fidelity solutions obtained with $N_{c}=10$ and $N_{c}=80$ checkpoints, respectively.
@@
Figure 21. Phase field isocontours for Rayleigh-Taylor instability at $t^{*}=2.95$ and $\mathrm{At}=0.35$. In each subfigure the left field is given by the HF solution. In each subfigure, on the right, are shown respectively: LF solution, corrected LF solution with $N_{c}=10$ and corrected LF solution with $N_{c}=80$.
We observe that even for the smallest value of $N_{c}$ considered here, the OT-based correction substantially improves the overall prediction. Increasing the number of checkpoints to $N_{c}=80$ further enhances the accuracy, to the point that the corrected low-fidelity solution becomes almost indistinguishable from the high-fidelity one.
@@
More importantly, for this specific problem, where breakup phenomena are widespread, we assess the capability of the MF-OT-ROM approach to accurately predict the total interfacial surface from a quantitative perspective. In Figure 22, we report the normalized interface surface as a function of time for both fidelities, considering the uncorrected and OT-corrected low-fidelity solutions. First, the temporal evolution clearly reflects the breakup dynamics. At early times, the interface between the two phases is relatively smooth and the total surface is limited. As the instability develops, gravitational effects enhance mixing, generating complex structures and significantly increasing the contact area between the phases.
@@
Figure 22. Normalized interfacial surface as a function of time for the Rayleigh-Taylor instability problem. High-fidelity solution is represented by the dash-dotted black line, uncorrected LF by the dashed black line and corrected LFs are shown using color gradient from blue to red for increasing values of $N_{c}$.
@@
Second, a classical resolution effect can be observed: coarse simulations are generally unable to capture small-scale features, such as vortices and thin interfacial ligaments. As a consequence, the low-fidelity solution systematically under-predicts the total interfacial area.
@@
Finally, as the number of checkpoints $N_{c}$ increases, the MF-OT-ROM prediction progressively converges toward the reference high-fidelity solution, confirming the effectiveness of the OT-based correction in recovering fine-scale interfacial dynamics. Interestingly, the MF-OT-ROM curve approaches the HF reference from above. This behavior may be attributed to the regularization employed within the OT solver. Indeed, as observed in the middle panel of Figure 21, a limited number of checkpoints produces a solution that appears less smooth. This residual lack of regularity can artificially increase the small-scale oscillations at the interface, ultimately leading to a slight over-prediction of the total interfacial area.
@@ 6.2.2. PMF-OT-ROM
In this last part we test the parametric multi-fidelity approach for the Rayleigh-Taylor Instability problem. It is worth noting that we expect significant variations in the dynamics for different parameters.
In Figure 23 we compare how the relative error behaves for different number of checkpoints. Normally, we would expect that for larger numbers of checkpoints the error would decrease. This is generally the case but in a different manner depending on the problem’s dynamics. In fact, at early stages, the improvement given by considering more checkpoints is significant. The initial dynamics, in fact, is very similar across the different parameters as the initial condition for the phase field is exactly the same across the whole dataset. However, as the dynamics of the problem unfolds, the type of breakups obtained at later stages is significantly different. At those stages, considering more checkpoints does not significantly improve the prediction of the OT solver. This clearly indicates that, as already observed previously, the benefit of using more checkpoints to reconstruct the fast-dynamical features of the problem is limited by the naive choice of the linear interpolator for the virtual time $\alpha$. The same concept in terms of mean error can be seen in Figure 24, where we can still observe that the error decrease by increasing the number of checkpoints used in the OT strategy, but at a slower rate.
Figure 23. Relative error between low-fidelity and high-fidelity for different number of checkpoints. Horizontal dashed lines represent the mean error for each temporal trajectory.
@@
Within this framework we also modified the total number of snapshots at disposal so that we can observe the same trends in terms of parametric refinements. In Figure 25 we compare how the relative error behaves for different number of checkpoints and different size of the nominal dataset. Normally, we would expect that given more data the error would decrease. In this scenario, at late stages of the problem’s dynamics, considering more data improves the prediction of the OT solver, even better when a large number of checkpoints is considered (left panel of Figure 25).
Figure 25. Relative error between low- and high-fidelity using the full dataset (circles) and half of it (squares). Dashed black line, uncorrected LF solution; dotted line, mean relative error for uncorrected LF solution. Left, $N_{c}=10$; right, $N_{c}=80$.
The same concept in terms of mean error is shown in Figure 26 where we can observe that indeed the error decrease both by enlarging the initial dataset and by increasing the number of checkpoints used in the OT solver. In general it is then interesting to notice that a certain tradeoff between temporal and parametric refinements is necessary in order to achieve an optimal result.
@@
As a more informative metric for this test case, we examine the total interface surface area integrated over the domain as a function of time. This quantity is reported in Figure 27, where we compare the predictions obtained with different numbers of checkpoints and different levels of parametric refinement.
In the early stages of the evolution, even a relatively small number of checkpoints ($N_{c}=10$) is sufficient to accurately capture the interface dynamics. However, as time progresses and increasingly complex breakup events occur, discrepancies between the OT prediction and the HF solution become more pronounced. This deviation reflects the growing complexity of the interface topology, which is more challenging to reconstruct with a limited number of trajectories.
@@
Figure 27. Normalized interface area predicted using the full dataset (circles) and half of it (squares). Left, $N_{c}=10$; right, $N_{c}=80$.
@@
In Figure 28 we compare the fields in the parametric setting as well. We can notice that for $N_{c}=10$ the differences are quite noticeable but if we increase the number of checkpoints to $N_{c}=80$ we obtain a quite good agreement between the high-fidelity solution and the OT-corrected low-fidelity one. Similarly to the predicted total interfacial area, we are obtaining a good prediction in the parametric setting too although not as accurate as the single parametric multi-fidelity approach presented earlier.
@@
Figure 28. Phase field isocontours for Rayleigh-Taylor instability at $t^{*}=2.95$ and $\mathrm{At}=0.35$. In each subfigure the left field is given by the HF solution. In each subfigure, on the right, are shown respectively: LF solution, corrected LF solution with $N_{c}=10$ and corrected LF solution with $N_{c}=80$.
In the previous sections, we presented several combinations of OT-based approaches (MF-OT-ROM and PMF-OT-ROM) and different test cases (Rider-Kothe vortex and Rayleigh-Taylor instability). In all these analyses, however, we considered only a single test parameter to validate the proposed approaches. In this final section, we focus on investigating how the most general approach (PMF-OT-ROM) performs in the most challenging scenario (Rayleigh-Taylor instability) across different parametric settings. In particular, for this problem, we expect increasingly complex dynamics as the Atwood number is raised. Previous analyses were performed at $\mathrm{At}=0.35$. In the present study, we additionally consider $\mathrm{At}=0.5$ and $\mathrm{At}=0.65$.
Figure 29 presents a qualitative comparison of high-fidelity versus corrected low-fidelity solutions for the different Atwood numbers. For relatively small Atwood numbers (left panel), the breakups are mild, with the high-fidelity solution predicting very thin ligaments and a small number of dispersed droplets. As the Atwood number increases, more complex interfacial patterns emerge, making it challenging for the reduced model to capture all fine-scale details accurately. Nevertheless, the large-scale dynamics, such as the position of the lower plume and the overall shape of the dominant interfacial structures, are correctly predicted.
Figure 29. Phase-field isocontours for the Rayleigh-Taylor instability at $t^{*}=2.95$ for different Atwood numbers. In each subfigure, the left field corresponds to the HF solution, and the right shows the corrected LF solution with $N_{c}=80$. From left to right: $\mathrm{At}=0.35,0.5,0.65$.
Figure 30 shows the relative error for two Atwood numbers. Although a large number of checkpoints is used, increasing the parametric sampling does not significantly improve the predictions for high Atwood numbers in the late stages of the Rayleigh-Taylor instability. This behavior is expected due to the highly complex breakup dynamics that occur at larger Atwood numbers.
Figure 30. Relative error between low- and high-fidelity solutions using the full dataset (circles) and half of it (squares) with $N_{c}=80$. Dashed black line: uncorrected LF solution; dotted black line: mean relative error of the uncorrected LF solution. Left, $\mathrm{At}=0.35$; right, $\mathrm{At}=0.65$.
In terms of predicted interfacial area, Figure 31 shows good agreement between the best-performing model ($N_{c}=80$ and $\Delta\mathrm{At}=0.025$) and the high-fidelity solution. For $\mathrm{At}=0.35$, the dynamics are accurately captured throughout the simulation. For $\mathrm{At}=0.65$, the interfacial area is well approximated up to $t^{*}=2.5$, after which complex breakup events lead to a deterioration of the reduced solution.
Figure 31. Normalized interfacial area predicted using the full dataset (circles) and half of it (squares) with $N_{c}=80$. Left, $\mathrm{At}=0.35$; right, $\mathrm{At}=0.65$.
@@ 7. Conclusions and Perspectives
In this paper, we have introduced two extensions to the Optimal Transport-based ROM framework. The multi-fidelity residual interpolation provides a method for correcting low-fidelity models by interpolating the HF-LF discrepancy field in time using displacement interpolation. The parametric displacement interpolation extends the framework to handle additional physical parameters beyond time, enabling prediction at unseen parameter values through a combination of parametric and temporal interpolation. For multi-fidelity parametric problems, we presented two strategies: interpolating the HF solution directly or interpolating the residual field in parameter space. This provides a computationally efficient and flexible approach for exploring parameter-dependent systems.
The mathematical formulation of these frameworks is general and can be applied to a wide range of problems governed by PDEs. By framing the discussion in the context of challenging diffuse-interface methods, we highlight the potential of these methods to handle complex, nonlinear dynamics involving moving fronts and interfaces.
Several directions for future work can be identified. First, the selection of training parameters and time checkpoints could be made adaptive, using error estimators to place computational effort where it is most needed. Second, the linear interpolation of the OT map in parameter space could be extended to barycentric interpolation for multiple parameters. Third, machine learning methods such as Gaussian Process Regression could be used to learn the mapping from physical parameters to the OT plans or interpolation weights. Finally, numerical validation on two-phase flow modelling problems, will quantify the performance and limitations of the proposed frameworks.

### 27 Mar 2026 v2

@@ 7.2. Rayleigh-Taylor instability
We now consider an even more complex benchmark represented by the Rayleigh-Taylor instability, which occurs when an interface between two fluids with different densities experiences a pressure gradient opposing the density gradient [87]. The computational domain is $d\times 4d]$ with $d=1$, and the interface is initially defined as the curve $y(x)=2d+0.1d\cos{(2\pi x/d)}$. The Rayleigh-Taylor instability is characterized by the Reynolds number $\mathrm{Re}=(\rho_{1}d^{3/2}\|\textbf{g}\|^{1/2})/\mu$ and the Atwood number $\mathrm{At}=(\rho_{1}-\rho_{2})/(\rho_{1}+\rho_{2})$ which are classically set to $3000$ and $0.5$. Time is adimensionalized as $t^{*}=t/\sqrt{d/(\|\textbf{g}\|\mathrm{At})}$, while for the EOS relation in Equation (26) we choose the following parameters $\gamma_{1}=\gamma_{2}=1.4$, $P^{\infty}_{1}=P^{\infty}_{2}=6000$.
The Reynolds number characterizes the relative importance of inertial to viscous effects in the coupled system and is governed by gravity and the mixture viscosity, which is assumed constant across the two phases for this case. The Atwood number, on the other hand, quantifies the density ratio between the phases. When $\mathrm{At}=0$, the two phases have identical densities, whereas increasing values of the Atwood number correspond to progressively larger density ratios. As the Atwood number increases, one therefore expects the flow to develop increasingly complex interfacial patterns.
The top boundary is treated as a Riemann-invariant boundary condition with zero velocity and constant pressure, the bottom boundary is a no-slip wall whereas slip-wall boundary conditions are prescribed on the other lateral sides of the domain. For this study, two different meshes with quadrilateral elements are considered, involving 4040k and 160160k degrees of freedom, LF and HF, respectively. A $4^{\mathrm{th}}$ order Runge-Kutta scheme was used for time integration. For this second test case, we consider a dataset consisting of $301$ snapshots, corresponding to a time window of $\Delta\tau=0.01$. The time step sizes used in the simulations are approximately $\Delta t_{40\mathrm{k}}\approx 10^{-5}$ and $\Delta t_{160\mathrm{k}}\approx 5\cdot 10^{-6}$ for the two spatial resolutions considered, to satisfy CFL constraints.
Concerning its parametric extension, we consider a variation of Atwood numbers in the following range:
@@
An overview of the parametric problem’s dynamics is shown in Figure 19. If not specified otherwise, we will always consider the $\mathrm{At}=0.35$ case to test our framework for the Rayleigh-Taylor instability.
@@
Figure 19. LF (top) and HF (bottom) phase-field isocontours at different Atwood numbers for the Rayleigh-Taylor instability at $t^{*}=3.0$.
@@ 7.2.1. MF-OT-ROM
A markedly different trend emerges compared to the previous test case: here, the dynamics accelerates over time. Initially, the interface evolves relatively slowly, but it subsequently accelerates due to gravitational effects, leading to complex breakup patterns. This behavior is also reflected in the MF-OT-ROM procedure, as the errors tend to grow as time progresses. This indicates that a more involved regression of the virtual time $\alpha$ may be needed to better capture the accelerating dynamics. Despite this, the current approximation already provides a significant improvement over the uncorrected LF solution, even with a limited amount of data, and increasing the number of checkpoints significantly improves the overall accuracy of the method, consistently with the behavior observed in the previous test case.
@@
In Figure 21, we present the predicted field of $\phi$ for the Rayleigh-Taylor instability problem. Each panel is split into two halves, where the HF solution is mirrored against the corresponding LF prediction (either corrected or uncorrected) to facilitate visual comparison. From left to right, the level of OT correction increases. The left panel displays the uncorrected LF solution, while the central and right panels show the corrected LF solutions obtained with $N_{c}=10$ and $N_{c}=80$ checkpoints, respectively. We observe that even for the smallest value of $N_{c}$ considered here, the OT-based correction substantially improves the overall prediction. Increasing the number of checkpoints to $N_{c}=80$ further enhances the accuracy, to the point that the corrected LF solution becomes almost indistinguishable from the HF one.
@@
Figure 21. Phase-field isocontours for Rayleigh-Taylor instability at $t^{*}=2.95$ and $\mathrm{At}=0.35$. In each subfigure, the left field is given by the HF solution, while right field shows respectively: LF solution, corrected LF solution with $N_{c}=10$ and corrected LF solution with $N_{c}=80$.
@@
More importantly, for this specific problem, where breakup phenomena are widespread, we assess the capability of the MF-OT-ROM approach to accurately predict the total interfacial surface from a quantitative perspective. In Figure 22, we report the normalized interface surface as a function of time for both fidelities, considering the uncorrected and OT-corrected LF solutions. First, the temporal evolution clearly reflects the breakup dynamics. At early times, the interface between the two phases is relatively smooth, and the total surface is limited. As the instability develops, gravitational effects enhance mixing, generating complex structures and significantly increasing the contact area between the phases.
@@
Figure 22. Normalized interfacial surface as a function of time for the Rayleigh-Taylor instability problem. Dash-dotted line, HF; dashed, uncorrected LF; color gradient, corrected LFs.
@@
Second, a classical resolution effect can be observed: coarse simulations are generally unable to capture small-scale features, such as vortices and thin interfacial ligaments. As a consequence, the LF solution systematically under-predicts the total interfacial area.
@@
Finally, as the number of checkpoints $N_{c}$ increases, the MF-OT-ROM prediction progressively converges to the reference HF solution, confirming the effectiveness of the OT-based correction in recovering fine-scale interfacial dynamics. Interestingly, the MF-OT-ROM curve approaches the HF reference from above. This behavior may be attributed to the regularization employed within the OT solver. Indeed, as observed in the middle panel of Figure 21, a limited number of checkpoints produces a solution that appears less smooth. This residual lack of regularity can artificially increase the small-scale oscillations at the interface, ultimately leading to a slight over-prediction of the total interfacial area.
@@ 7.2.2. PMF-OT-ROM
Finally, we test the parametric multi-fidelity approach for the Rayleigh-Taylor instability problem, for which we have observed significant variations in the dynamics for different parameters in Figure 19.
In Figure 23 we compare how the relative error behaves for different number of checkpoints. Normally, we would expect that for larger numbers of checkpoints the error would decrease. This is generally the case but in a different manner depending on the problem’s dynamics and its “velocity”. At early stages, the improvement given by considering more checkpoints is significant. In fact, the initial dynamics is very similar across the different parameters as the initial condition for the phase-field is exactly the same across the whole dataset. However, as the dynamics of the problem unfolds, the type of breakups obtained at later stages is significantly different. At those stages, considering more checkpoints in time does not significantly improve the prediction of the OT solver. This clearly indicates that, as already previously observed, the benefit of using more checkpoints to reconstruct the fast-dynamical features of the problem is limited by the naive choice of the linear interpolator for the virtual time $\alpha$. The same concept in terms of mean error can be seen in Figure 24, where we can still observe that the error decrease by increasing the number of checkpoints used in the OT strategy, but at a slower rate.
Figure 23. Relative error between LF and HF for different number of checkpoints. Horizontal dashed lines represent the mean error for each temporal trajectory.
@@
Within this framework we also modified the total number of snapshots at disposal so that we can observe the same trends in terms of parametric refinements. In Figure 25 we compare how the relative error behaves for different number of checkpoints and different number of parametric trajectories considered. Normally, we would expect that given more parametric trajectories the error would decrease. In this scenario, at late stages of the problem’s dynamics, considering more data improves the prediction of the overall procedure, even better when a large number of temporal checkpoints is considered (left panel of Figure 25).
Figure 25. Relative error between low- and high-fidelity using the full dataset (circles) and half of it (squares), for $N_{c}=10$ and $N_{c}=80$, left and right respectively. Dashed black line, uncorrected LF solution; dotted line, mean relative error for uncorrected LF solution.
The same concept in terms of mean error is shown in Figure 26 where we can observe that indeed the error decreases both by enlarging the initial dataset or alternatively by increasing the number of checkpoints used in the OT solver. In general, it is then interesting to notice that a certain trade-off between temporal and parametric refinements is necessary in order to achieve an optimal result.
@@
As a more informative metric for this test case, we examine the total interfacial area as a function of time. This quantity is reported in Figure 27, where we compare the predictions obtained with different numbers of checkpoints and different levels of parametric refinement.
In the early stages of the evolution, even a relatively small number of checkpoints ($N_{c}=10$) is sufficient to accurately capture the interfacial dynamics. However, as time progresses and increasingly complex breakup events occur, discrepancies between the OT prediction and the HF solution become more pronounced. This deviation reflects the growing complexity of the interfacial topology, which is more challenging to reconstruct with a limited number of trajectories.
@@
Figure 27. Normalized interface area predicted using the full dataset (circles) and half of it (squares) for $N_{c}=10$ and $N_{c}=80$, left and right respectively.
@@
In Figure 28 we compare the fields in the parametric setting as well. We can notice that for $N_{c}=10$ the differences are quite noticeable but if we increase the number of checkpoints to $N_{c}=80$ we obtain a quite good agreement between the HF solution and the OT-corrected LF one. Similarly to the predicted total interfacial area, we are obtaining a good prediction in the parametric setting too although not as accurate as the single parametric multi-fidelity approach presented earlier.
@@
Figure 28. Phase-field isocontours for Rayleigh-Taylor instability at $t^{*}=2.95$ and $\mathrm{At}=0.35$. In each subfigure, the left field is given by the HF solution, while right field shows respectively: LF solution, corrected LF solution with $N_{c}=10$ and corrected LF solution with $N_{c}=80$.
In the previous analysis, we tested OT-based approaches on a single test parameter to validate the proposed methodologies. Thus, we conclude the investigation by examining the performance of the most general approach, namely the PMF-OT-ROM method, in the most challenging scenario considered in this work, i.e. the Rayleigh-Taylor instability, across the entire range of parametric configurations.
Figure 29 reports the error analysis performed over the entire parametric range in such setting, from which a few relevant trends can be observed. In general, the error tends to increase as the Atwood number grows, up to approximately $\mathrm{At}=0.5$. This behavior is expected, since for relatively small Atwood numbers the overall dynamics of the system remains more regular and less affected by strong interfacial breakups. Conversely, larger Atwood numbers lead to increasingly complex interfacial dynamics, characterized by stronger nonlinear effects and fragmentation phenomena (see Figure 19).
Figure 29. Mean relative error analysis across the full parametric range with $\Delta\mathrm{At}=0.05$ (top) and $\Delta\mathrm{At}=0.025$ (bottom).
Concerning the PMF-OT-ROM approach, we observe that, even with a relatively small number of checkpoints ($N_{c}=10$), the corrected solution consistently improves the baseline uncorrected LF one for all the parametric trajectories. Furthermore, in agreement with previous observations for the reference parametric configuration, increasing the number of checkpoints significantly improves the accuracy of the corrected solution. This improvement appears to saturate relatively quickly when considering the coarser parametric spacing ($\Delta\mathrm{At}=0.05$). In contrast, for the finer parametric resolution ($\Delta\mathrm{At}=0.025$), a noticeable improvement in accuracy is observed when increasing the number of checkpoints from $N_{c}=20$ to $N_{c}=80$. This gain in accuracy is particularly evident for lower Atwood numbers, where the flow dynamics remain comparatively less complex.
@@ 8. Conclusions and Perspectives
In this paper, we have introduced two extensions to the Optimal Transport-based ROM framework. The multi-fidelity residual interpolation provides a method for correcting low-fidelity models by interpolating the HF-LF discrepancy field in time using displacement interpolation (MF-OT-ROM). The parametric displacement interpolation extends the framework to handle additional physical parameters beyond time, enabling prediction at unseen parameter values through a combination of parametric and temporal interpolation (PMF-OT-ROM). For multi-fidelity parametric problems, the HF solution is first interpolated across neighboring trajectories to produce synthetic data. These data are then used to compute residuals, which are subsequently employed to correct the LF solutions for previously unseen parametric configurations.
The mathematical formulation of these frameworks is general and can be applied to a wide range of problems governed by PDEs. By framing the discussion in the context of challenging diffuse-interface methods, we highlight the potential of these methods to handle complex, nonlinear dynamics involving moving fronts and interfaces. We have successfully applied the proposed extensions to two standard test cases for two-phase flow modeling, namely the Rider-Kothe vortex and the Rayleigh-Taylor instability. Both problems are characterized by unsteady dynamics with sharp structures that evolve and propagate in space. The Rayleigh-Taylor instability, in particular, involves complex interfacial dynamics with frequent breakup events, which become even more pronounced at large Atwood numbers. In both test cases, the proposed extensions significantly improve the baseline predictions, highlighting the potential of OT-based strategies for surrogate modeling in two-phase flows.
Several directions for future work can be identified. First, the selection of training parameters and time checkpoints could be made adaptive, using error estimators to place computational effort where it is most needed. Second, the linear interpolation of the OT map in parameter space could be extended to barycentric interpolation for multiple parameters. Third, machine learning methods such as Gaussian Process Regression could be used to learn the mapping from physical parameters to the OT plans or interpolation weights. Finally, OT methodologies could be further leveraged beyond surrogate modeling, for instance in the modeling of turbulent two-phase flows, such as predicting breakup dynamics.